OpenStudy (anonymous):

Solve the equation. Be sure to show your work. All final answers must be in simplest form. Student Answer: x-7/9 = 2/27 = x- 7/9+ 7/9 = 2/27 + 7/9 =x= 9/36 x= 1/3 check: 9/36 - 7/9 = 2/27 Thus the answer is x= 1/3 simplest form. What did I do wrong?

OpenStudy (anonymous):

idk bcoz i m in 7th grade

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

give me medal

OpenStudy (anonymous):

\[\frac{2}{27} + \frac{7}{9} \neq \frac{9}{36}\]

OpenStudy (anonymous):

2/27+7/9≠9/36

OpenStudy (anonymous):

0k. I see. I should of multiplied 7/9* 3

OpenStudy (anonymous):

x - 7/9 = 2/27 x - 21/27 = 2/27 x = 2/27 + 21/27 x = 23/27

OpenStudy (anonymous):

ok. see where I went wrong. thanks.

OpenStudy (anonymous):

best way is to convert fractions to sane deniminator 7/9 = 7*3 / 9*3 = 21/27

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I got it. then mutiple to get the answer. thanks for the help.

OpenStudy (anonymous):

ok